# Revisiting Non-Acyclic GFlowNets in Discrete Environments

### v1

@@ 3.5 Training Non-Acyclic GFlowNets
$\mathcal{F}(x\to s_{f})=\mathcal{R}(x)\;\;\forall(x\to s_{f})\in\mathcal{E}\,,$ (9)
@@
An important fact from the literature on acyclic GFlowNets (Malkin et al., 2022; Bengio et al., 2023) that was overlooked in the work of (Brunswic et al., 2024), but holds in non-acyclic case as well, is that it is generally easy to manually pick a backward policy such that the induced trajectory distribution will satisfy reward matching condition. A simple and natural choice is to take $\mathcal{P}_{\mathrm{B}}(x\mid s_{f})=\mathcal{R}(x)/\mathcal{Z}$ for $s_{f}$ and fix $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=1/|\operatorname{in}(s^{\prime})|$ for all other states. It is worth mentioning that $\mathcal{Z}$ is generally unknown, but this issue is circumvented in GFlowNets by learning unnormalized flows or making $\mathcal{Z}$ itself a learnable parameter depending on the chosen loss function (Malkin et al., 2022; Bengio et al., 2023; Madan et al., 2023). Moreover, Proposition 3.8 shows the uniqueness of the corresponding $\mathcal{P}_{\mathrm{F}}$. Thus, we state the main practical corollary of this result:
@@
The main disadvantage of learning with a fixed backward policy in non-acyclic GFlowNets that does not arise in acyclic GFlowNets is the fact that expected trajectory length $\mathbb{E}_{\tau\sim\mathcal{P}}[n_{\tau}]$ of a manually chosen $\mathcal{P}_{\mathrm{B}}$ can be large. A natural way to circumvent this issue is to consider a learnable backward policy, which is also a widely employed choice in acyclic GFlowNet literature (Malkin et al., 2022; Jang et al., 2024; Gritsaev et al., 2024). However, (Brunswic et al., 2024) made an important discovery by pointing out that standard losses from acyclic GFlowNet literature are not stable (Theorem 3), meaning that the expected trajectory length can grow uncontrollably during training. The concept of stability was introduced with respect to learnable edge flows (Definition 3), which implies that the corresponding backward policy also changes during training. Using a stable loss, e.g. (5), was proposed as way to approach this issue. At the same time, we argue that efficient training of a non-acyclic GFlowNet with controlled expected trajectory length in case of a learnable $\mathcal{P}_{\mathrm{B}}$ is possible without utilizing stable losses. The next proposition is a simple corollary of Definition 3.5:
@@
We also believe that exploiting this equivalence is a crucial direction for future research on non-acyclic GFlowNets. We further explore a particular solution, which suggests the use of a state flow as a regularizer in the existing GFlowNet losses. We consider an example of the detailed balance loss DB (4). In this case, Proposition 3.9 implies that learning a non-acyclic GFlowNet with the smallest expected trajectory length can be formulated as the following constraint optimization problem:
@@
Finally, it is important to mention that flow-based regularizers in a non-acyclic case were already proposed in Theorem 1 of (Brunswic et al., 2024), but only for the stable loss setup. Moreover, they were introduced in order to find an acyclic flow. Our paper further explores and sheds new light on this phenomenon, showing that training can be carried out even when an unstable loss is utilized with regularization. Moreover, when the total flow is minimized, one can ensure the smallest possible expected trajectory length.
@@ 3.6 Connections with Entropy-Regularized RL
Let $\mathcal{G}$ be a graph of a non-acyclic GFlowNet, $\mathcal{R}$ a GFlowNet reward, and $\mathcal{P}_{\mathrm{B}}>0$ a fixed backward policy that satisfies the reward matching condition. Let $\mathcal{F}$ be the flow induced by $\mathcal{P}_{\mathrm{B}}$ with $\mathcal{F}(s_{f})=\mathcal{Z}$, and $\mathcal{P}_{\mathrm{F}}$ be a unique forward policy corresponding to $\mathcal{P}_{\mathrm{B}}$ (see Proposition 3.8). Define a deterministic MDP $\mathcal{M}_{\mathcal{G}}$ induced by a graph $\mathcal{G}$, where the state space $\mathcal{S}$ coincide with vertices of $\mathcal{G}$, action space $\mathcal{A}_{s}$ for each state $s$ correspond to $\operatorname{out}(s)$, and transition kernel is defined as transition in the graph $\text{P}(s^{\prime}\mid s,a)=\mathbb{I}\{a=s^{\prime}\}$, $a\in\operatorname{out}(s)$. We use no discounting ($\gamma=1.0$) and set RL rewards for terminating transitions $r(x,s_{f})=\log\mathcal{R}(x)$, and for all other transitions $r(s,s^{\prime})=\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Then, the following statement holds.
@@
The optimal policy $\pi^{\star}_{\lambda=1}(s^{\prime}\mid s)$ for the entropy-regularized MDP $\mathcal{M}$ with coefficient $\lambda=1$ is equal to $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$ for all $s\in\mathcal{S}\setminus\{s_{f}\},s^{\prime}\in\mathcal{A}_{s}$. Moreover, regularized optimal value $V^{\star}_{\lambda=1}(s)$ and Q-value $Q^{\star}_{\lambda=1}(s,s^{\prime})$ coincide with $\log\mathcal{F}(s)$ and $\log\mathcal{F}(s\to s^{\prime})$ respectively.
@@ 4 Experiments
Scaling hypothesis. When $\mathcal{P}_{\mathrm{B}}$ is trainable, the main factor that plays a crucial role in loss stability in practice, i.e., controlled mean trajectory length of the trained non-acyclic GFlowNet, is the scale in which the error between flows is computed. Indeed, the standard DB loss (4) operates in log-flow scale $\Delta\log\mathcal{F}$, while standard SDB (5) operates in flow scale $\Delta\mathcal{F}$. We hypothesize that using log-flow scale losses without regularization can lead to arbitrary large trajectory length, while flow scale losses are biased towards solutions with smaller flows and thus do not suffer from this issue.
In this section, we use DB or SDB to specify the utilized loss, $\Delta\log\mathcal{F}$ or $\Delta\mathcal{F}$ to specify the flow scale used to compute the error, and use $\lambda=C$ to specify the strength of the proposed state flow regularization (see Section 3.5). For example, $(\texttt{DB},\Delta\log\mathcal{F})$ in the legend correspond to (4), $(\texttt{SDB},\Delta\mathcal{F})$ corresponds to (5) and $(\texttt{DB},\Delta\log\mathcal{F},\lambda=C)$ corresponds to (12). Detailed discussion on loss scaling and stability is provided in Section B.2.
@@ 4.1 Experimental Setting
In both environments, we consider two settings: training with a fixed backward policy $\mathcal{P}_{\mathrm{B}}$ that is almost uniform and using a trainable $\mathcal{P}_{\mathrm{B}}$. In the second case, the initial log flow $\log\mathcal{F}_{\theta}(s_{0})=\log\mathcal{Z}_{\theta}$ is also being learned. Thus, we can examine its convergence to the true normalizing constant $\log\mathcal{Z}$. See Appendix B.3 for details.
4.2 Hypergrids
We start with non-acyclic hypergrid environments (Brunswic et al., 2024). These environments are small enough that the normalizing constant $\mathcal{Z}$ can be computed exactly, and the trained sampler can be efficiently evaluated against the exact reward distribution. States are points with integer coordinates $s\in\{0,\ldots,H-1\}^{D}$ inside a $D$-dimensional hypercube with side length $H$, plus two auxiliary states $s_{0}$ and $s_{f}$. Possible transitions correspond to increasing or decreasing any coordinate by $1$ without exiting the grid. Moreover, each state has a terminating transition $s\to s_{f}$, thus $\mathcal{X}=\mathcal{S}\setminus\{s_{0},s_{f}\}$. The reward function has modes near the grid corners, separated by wide troughs with very small rewards. To measure sampling accuracy, total variation distance is computed between the true reward distribution $\mathcal{R}(x)/\mathcal{Z}$ and an empirical distribution of the last $2\cdot 10^{5}$ samples seen in training, which coincides with $\frac{1}{2}$ of $L^{1}$ distance on discrete domains.
@@ 4.2 Hypergrids
Table 1: Comparison on the permutation environment. $C(k)\;L^{1}$ is $L^{1}$ distance between true and empirical distribution of fixed point probabilities $C(k)$, $\Delta\mathcal{R}$ is relative error of mean reward proposed in (Shen et al., 2023), $\Delta\log\mathcal{Z}$ is $|\log\mathcal{Z}_{\theta}-\log\mathcal{Z}|$. Mean and std values are computed over 3 random seeds. Blue indicates the best metric, red indicates the smallest expected trajectory length.
@@
We begin our analysis with a $7\times 7$ grid to study the effects of learning under a fixed backward policy compared to a trainable backward policy. Since the environment is small, it is possible to find the flows induced by the fixed $\mathcal{P}_{\mathrm{B}}$ exactly, thus also its expected trajectory length, see Appendix B.4. Figure 1 presents the results. First, we note that both $(\texttt{DB},\Delta\log\mathcal{F})$ and $(\texttt{SDB},\Delta\mathcal{F})$ with fixed $\mathcal{P}_{\mathrm{B}}$ converge to the true expected trajectory length induced by the fixed backward policy, which is in line with Corollary 3.11. However, for all losses, using trainable $\mathcal{P}_{\mathrm{B}}$ allows us to find a solution with a smaller trajectory length. In addition, we observe that using a loss in $\Delta\mathcal{F}$ scale results in slower convergence and a slight bias in the learned forward policy than in case of $\Delta\log\mathcal{F}$ scale, for both fixed and learned $\mathcal{P}_{\mathrm{B}}$. Finally, an interesting note is that using an unstable DB loss in $\Delta\log\mathcal{F}$ scale without state flow regularization can still result in a small expected trajectory length, as we see in this experiment. However, we further show that this is not the case for a larger environment.
@@ 4.3 Permutations
Next, we consider the environment corresponding to the Cayley Graph of the group of permutations on $n$ elements $\{1,2,\dots,n\}$ from (Brunswic et al., 2024). Each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ is a permutation of fixed length $(s(1),\dots,s(n))$, and there are $n-1$ possible transitions that correspond to swapping a pair of adjacent elements $s(k)$ and $s(k+1)$, plus a transition that corresponds to a circular shift of the permutation to the right $(s(n),s(1),\dots,s(n-1))$. In addition, each state has a terminating transition $s\to s_{f}$. GFlowNet reward utilized in the experiments of (Brunswic et al., 2024) is $\mathbb{I}[s(1)=1]$. We argue that this results in a fairly simple task, and a trivial forward policy exists that just applies circular shift until $1$ is on the first position. We opt for using a more complex reward distribution in our experiments and define GFlowNet reward in terms of the number of fixed points in a permutation $\mathcal{R}(s)=\exp\mathopen{}\mathclose{{}\left(\frac{1}{2}\sum_{k=1}^{n}\mathbb{I}\{s(k)=k\}}\right)$.
Since with the growth of $n$, the number of states $n!$ quickly becomes too large to compute total variation distance as it was done for hypergrid, we track convergence of a number of statistics to their true respective values. Firstly, we compute the relative error between mean reward of GFlowNet samples and true expected reward as it was proposed in (Shen et al., 2023). Secondly, denote $C(k)$ as the probability that a permutation sampled from the reward distribution has $k$ fixed points. We compute $L^{1}$ error between the vector $(C(0),C(1),\dots,C(n))$ and its empirical estimate over last $10^{5}$ samples seen in training. Finally, we track the convergence of the trained $\log\mathcal{Z}_{\theta}$ to the true value of $\log\mathcal{Z}$. In Appendix C.3.1, we show how true reference values of these quantities can be efficiently computed.
@@ 5 Conclusion
Future work could explore applying other losses from acyclic GFlowNet literature (Madan et al., 2023; da Silva et al., 2024; Hu et al., 2024) to the non-acyclic setting. Based on Theorem 3.13, another promising direction is to apply RL techniques and algorithms to GFlowNets in the non-acyclic case, following their success for acyclic GFlowNets (Tiapkin et al., 2024; Mohammadpour et al., 2024; Lau et al., 2024; Morozov et al., 2024). Finally, environments where all states are terminal, i.e., have a transition into $s_{f}$, naturally arise in the non-acyclic case. Then, special modifications can be introduced to improve the propagation of a reward signal during training (Deleu et al., 2022; Pan et al., 2023).

### v2

@@ 3.5 Training Non-Acyclic GFlowNets
$\mathcal{F}(x\to s_{f})=\mathcal{R}(x)\;\;\forall x\to s_{f}\in\mathcal{E}\,,$ (9)
@@
An important fact from the literature on acyclic GFlowNets (Malkin et al., 2022; Bengio et al., 2023) that was overlooked in the work of (Brunswic et al., 2024), but holds in the non-acyclic case as well, is that it is generally easy to manually pick a backward policy such that the induced trajectory distribution will satisfy the reward matching condition. A simple and natural choice is to take $\mathcal{P}_{\mathrm{B}}(x\mid s_{f})=\mathcal{R}(x)/\mathcal{Z}$ for $s_{f}$ and fix $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=1/|\operatorname{in}(s^{\prime})|$ for all other states. It is worth mentioning that $\mathcal{Z}$ is generally unknown, but this issue is circumvented in GFlowNets by learning unnormalized flows or making $\mathcal{Z}$ itself a learnable parameter depending on the chosen loss function (Malkin et al., 2022; Bengio et al., 2023; Madan et al., 2023). Moreover, Proposition 3.8 shows the uniqueness of the corresponding $\mathcal{P}_{\mathrm{F}}$. Thus, we state the main practical corollary of this result:
@@
The main disadvantage of learning with a fixed backward policy in non-acyclic GFlowNets that does not arise in acyclic GFlowNets is the fact that the expected trajectory length $\mathbb{E}_{\tau\sim\mathcal{P}}[n_{\tau}]$ of a manually chosen $\mathcal{P}_{\mathrm{B}}$ can be large. A natural way to circumvent this issue is to consider a learnable backward policy, which is also a widely employed choice in acyclic GFlowNet literature (Malkin et al., 2022; Jang et al., 2024a; Gritsaev et al., 2025). However, (Brunswic et al., 2024) made an important discovery by pointing out that standard losses from acyclic GFlowNet literature are not stable (Theorem 3), meaning that the expected trajectory length can grow uncontrollably during training. The concept of stability was introduced with respect to learnable edge flows (Definition 3), which implies that the corresponding backward policy also changes during training. Using a stable loss, e.g. (5), was proposed as a way to approach this issue. At the same time, we argue that efficient training of a non-acyclic GFlowNet with controlled expected trajectory length in case of a learnable $\mathcal{P}_{\mathrm{B}}$ is possible without utilizing stable losses. The next proposition is a simple corollary of Definition 3.5:
@@
We also believe that exploiting this equivalence is a crucial direction for future research on non-acyclic GFlowNets. We further explore a particular solution, which suggests the use of a state flow as a regularizer in the existing GFlowNet losses. We consider an example of the detailed balance loss DB (4). In this case, Proposition 3.9 implies that learning a non-acyclic GFlowNet with the smallest expected trajectory length can be formulated as the following constrained optimization problem:
@@
Finally, it is important to mention that flow-based regularizers in the non-acyclic case were already proposed in Theorem 1 of (Brunswic et al., 2024), but only for the stable loss setup. Moreover, they were introduced in order to find an acyclic flow. Our paper further explores and sheds new light on this phenomenon, showing that training can be carried out even when an unstable loss is utilized with regularization. Moreover, when the total flow is minimized, one can ensure the smallest possible expected trajectory length. It is also worth pointing out that the idea of introducing a constrained optimization problem to accommodate for cycles in GFlowNets was mentioned in (Deleu, 2025).
@@ 3.6 Connections with Entropy-Regularized RL
Let $\mathcal{G}$ be a graph of a non-acyclic GFlowNet, $\mathcal{R}$ a GFlowNet reward, and $\mathcal{P}_{\mathrm{B}}>0$ a fixed backward policy that satisfies the reward matching condition. Let $\mathcal{F}$ be the flow induced by $\mathcal{P}_{\mathrm{B}}$ with $\mathcal{F}(s_{f})=\mathcal{Z}$, and $\mathcal{P}_{\mathrm{F}}$ be a unique forward policy corresponding to $\mathcal{P}_{\mathrm{B}}$ (see Proposition 3.8). Define a deterministic MDP $\mathcal{M}_{\mathcal{G}}$ induced by a graph $\mathcal{G}$, where the state space $\mathcal{S}$ coincides with vertices of $\mathcal{G}$, the action space $\mathcal{A}_{s}$ for each state $s$ corresponds to $\operatorname{out}(s)$, and the transition kernel is defined as transition in the graph $\text{P}(s^{\prime}\mid s,a)=\mathbb{I}\{a=s^{\prime}\}$, $a\in\operatorname{out}(s)$. We use no discounting ($\gamma=1.0$) and set RL rewards for terminating transitions $r(x,s_{f})=\log\mathcal{R}(x)$, and for all other transitions $r(s,s^{\prime})=\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Then, the following statement holds.
@@
The optimal policy $\pi^{\star}_{\lambda=1}(s^{\prime}\mid s)$ for the entropy-regularized MDP $\mathcal{M}_{\mathcal{G}}$ with coefficient $\lambda=1$ is equal to $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$ for all $s\in\mathcal{S}\setminus\{s_{f}\},s^{\prime}\in\mathcal{A}_{s}$. Moreover, regularized optimal value $V^{\star}_{\lambda=1}(s)$ and Q-value $Q^{\star}_{\lambda=1}(s,s^{\prime})$ coincide with $\log\mathcal{F}(s)$ and $\log\mathcal{F}(s\to s^{\prime})$ respectively for all $s\to s^{\prime}\in\mathcal{E}$.
@@ 4 Experiments
Scaling hypothesis. When $\mathcal{P}_{\mathrm{B}}$ is trainable, the main factor that plays a crucial role in loss stability in practice, i.e., controlled mean trajectory length of the trained non-acyclic GFlowNet, is the scale in which the error between flows is computed. Indeed, the standard DB loss (4) operates in log-flow scale $\Delta\log\mathcal{F}$, while standard SDB (5) operates in flow scale $\Delta\mathcal{F}$. We hypothesize that using log-flow scale losses without regularization can lead to arbitrarily large trajectory length, while flow scale losses are biased towards solutions with smaller flows and thus do not suffer from this issue.
In this section, we use DB or SDB to specify the utilized loss, $\Delta\log\mathcal{F}$ or $\Delta\mathcal{F}$ to specify the flow scale used to compute the error, and use $\lambda=C$ to specify the strength of the proposed state flow regularization (see Section 3.5). For example, $(\texttt{DB},\Delta\log\mathcal{F})$ in the legend corresponds to (4), $(\texttt{SDB},\Delta\mathcal{F})$ corresponds to (5) and $(\texttt{DB},\Delta\log\mathcal{F},\lambda=C)$ corresponds to (12). Detailed discussion on loss scaling and stability is provided in Section B.2.
@@ 4.1 Experimental Setting
Table 1: Comparison on the permutation environment. $C(k)\;L^{1}$ is the $L^{1}$ distance between the true and empirical distribution of fixed point probabilities $C(k)$, $\Delta\mathcal{R}$ is the relative error of mean reward proposed in (Shen et al., 2023), $\Delta\log\mathcal{Z}$ is $|\log\mathcal{Z}_{\theta}-\log\mathcal{Z}|$. Mean and std values are computed over 3 random seeds. Blue indicates the best metric, red indicates the smallest expected trajectory length.
@@
In both environments, we consider two settings: training with a fixed backward policy $\mathcal{P}_{\mathrm{B}}$ that is almost uniform and using a trainable $\mathcal{P}_{\mathrm{B}}$. In the second case, the initial log flow $\log\mathcal{F}_{\theta}(s_{0})=\log\mathcal{Z}_{\theta}$ is also being learned. Thus, we can examine its convergence to the logarithm of the true normalizing constant $\log\mathcal{Z}$. See Appendix B.3 for details.
4.2 Hypergrids
We start with non-acyclic hypergrid environments (Brunswic et al., 2024). These environments are small enough that the normalizing constant $\mathcal{Z}$ can be computed exactly, and the trained sampler can be efficiently evaluated against the exact reward distribution. States are points with integer coordinates $s\in\{0,\ldots,H-1\}^{D}$ inside a $D$-dimensional hypercube with side length $H$, plus two auxiliary states $s_{0}$ and $s_{f}$. Possible transitions correspond to increasing or decreasing any coordinate by $1$ without exiting the grid. Moreover, each state has a terminating transition $s\to s_{f}$, thus $\mathcal{X}=\mathcal{S}\setminus\{s_{0},s_{f}\}$. The reward function has modes near the grid corners, separated by wide troughs with very small rewards. To measure sampling accuracy, total variation distance is computed between the true reward distribution $\mathcal{R}(x)/\mathcal{Z}$ and an empirical distribution of the last $2\cdot 10^{5}$ samples seen in training, which coincides with $\frac{1}{2}$ of the $L^{1}$ distance on discrete domains.
We begin our analysis with a $7\times 7$ grid to study the effects of learning under a fixed backward policy compared to a trainable backward policy. Since the environment is small, it is possible to find the flows induced by the fixed $\mathcal{P}_{\mathrm{B}}$ exactly, thus also its expected trajectory length, see Appendix B.4. Figure 1 presents the results. First, we note that both $(\texttt{DB},\Delta\log\mathcal{F})$ and $(\texttt{SDB},\Delta\mathcal{F})$ with fixed $\mathcal{P}_{\mathrm{B}}$ converge to the true expected trajectory length induced by the fixed backward policy, which is in line with Corollary 3.11. However, for all losses, using trainable $\mathcal{P}_{\mathrm{B}}$ allows us to find a solution with a smaller trajectory length. In addition, we observe that using a loss in $\Delta\mathcal{F}$ scale results in slower convergence and a slight bias in the learned forward policy than in the case of $\Delta\log\mathcal{F}$ scale, for both fixed and learned $\mathcal{P}_{\mathrm{B}}$. Finally, an interesting note is that using an unstable DB loss in $\Delta\log\mathcal{F}$ scale without state flow regularization can still result in a small expected trajectory length, as we see in this experiment. However, we further show that this is not the case for a larger environment.
@@ 4.3 Permutations
Next, we consider the environment corresponding to the Cayley Graph of the of the symmetric group $\mathrm{S}_{n}$ (group of permutations on $n$ elements $\{1,2,\dots,n\}$) from (Brunswic et al., 2024). Each state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$ is a permutation of fixed length $(s(1),\dots,s(n))$, and there are $n-1$ possible transitions that correspond to swapping a pair of adjacent elements $s(k)$ and $s(k+1)$, plus a transition that corresponds to a circular shift of the permutation to the right $(s(n),s(1),\dots,s(n-1))$. In addition, each state has a terminating transition $s\to s_{f}$. GFlowNet reward utilized in the experiments of (Brunswic et al., 2024) is $\mathbb{I}[s(1)=1]$. We argue that this results in a fairly simple task, and a trivial forward policy exists that just applies circular shift until $1$ is in the first position. We opt for using a more complex reward distribution in our experiments and define GFlowNet reward in terms of the number of fixed points in a permutation $\mathcal{R}(s)=\exp\mathopen{}\mathclose{{}\left(\frac{1}{2}\sum_{k=1}^{n}\mathbb{I}\{s(k)=k\}}\right)$.
Since with the growth of $n$, the number of states $n!$ quickly becomes too large to compute total variation distance as it was done for hypergrid, we track convergence of a number of statistics to their true respective values. Firstly, we compute the relative error between the mean reward of GFlowNet samples and the true expected reward as it was proposed in (Shen et al., 2023). Secondly, denote $C(k)$ as the probability that a permutation sampled from the reward distribution has $k$ fixed points. We compute the $L^{1}$ error between the vector $(C(0),C(1),\dots,C(n))$ and its empirical estimate over the last $10^{5}$ samples seen in training. Finally, we track the convergence of the trained $\log\mathcal{Z}_{\theta}$ to the true value of $\log\mathcal{Z}$. In Appendix C.3.1, we show how true reference values of these quantities can be efficiently computed.
@@ 5 Conclusion
Future work could explore applying other losses from acyclic GFlowNet literature (Madan et al., 2023; Silva et al., 2024; Hu et al., 2025) to the non-acyclic setting. Based on Theorem 3.13, another promising direction is to apply known RL techniques and algorithms to GFlowNets in the non-acyclic case, following their success for acyclic GFlowNets (Tiapkin et al., 2024; Mohammadpour et al., 2024; Lau et al., 2024; Morozov et al., 2024). Finally, environments where all states are terminal, i.e., have a transition into $s_{f}$, naturally arise in the non-acyclic case. Then, special modifications can be introduced to improve the propagation of the reward signal during training (Deleu et al., 2022; Pan et al., 2023; Jang et al., 2024b).